# publications.bib

### v3

@@ 4.3. Verification of additional assumptions
Since $\delta_{B}$ is just the restriction of the character for the Borel of $\prod\mathrm{GSp}_{4}$ from the torus $\prod T$ to $T^{\prime}$, we see that the modulus character is
$\delta_{B}((t_{\tau 1},t_{\tau 2})_{\tau},t)=|t|^{-3d}\prod_{\tau}|t_{\tau 1}|^{4}|t_{\tau 2}|^{2}.$
Since $W_{G}=\prod_{\tau}W_{\mathrm{GSp}_{4}}$, consider any element $w=(w_{\tau})\in W_{G}$, where $w_{\tau}\in W_{\mathrm{GSp}_{4}}$. Observe that the expression obtained from the isomorphism (16) for $w=(w_{\tau})$ is simply the product of the isomorphisms for $\mathrm{GSp}_{4}$ each $w_{\tau}$. Thus, if we wanted to argue by contradiction, using the notation of the proof above, when $w_{\tau}=w_{i}$ for $i=1,\dots,7$, we should substitute for $t_{\tau 1},t_{\tau 2},t$ the values we considered above, subject to the additional constraint that we must have $t$, the similitude factor, being equal for all $\tau$.
We thus have two possibilities: either some $w_{\tau}$ is the Weyl group element $w_{2}$ (i.e. corresponding to the reflection
$w_{2}:a(t_{1},t_{2},t)\mapsto a(t_{1},t_{2}^{-1}t,t)$
or none of the $w_{\tau}$ are this element.
In the first situation, suppose that for some $\tau$, $w_{\tau}$ is the Weyl reflection $w_{2}$. If we consider the equation (16), evaluated on the element $t_{\tau 1}=x,t_{\tau 2}=x^{2}$,and $t_{\tau^{\prime}1}=t_{\tau^{\prime}2}=x^{2},t=x^{4}$ for all $\tau^{\prime}\neq\tau$, then equation (16) simplifies to
(19) $\mathbf{1}\simeq|x|^{-4},$
since one can check that substituting $t_{1}=t_{2}=x^{2},t=x^{4}$ into the isomorphism (16) for $\mathrm{GSp}_{4}$ for all the Weyl elements not equal to $w_{2}$ above simply gives the isomorphism $\mathbf{1}\simeq\mathbf{1}$ after simplification, and thus does not matter when taking products). This contradicts the banality assumption that $(p^{4}-1,\ell)=1$.
Now, we suppose we are in the second situation, i.e. no $w_{\tau}=w_{2}$. For any $w=(w_{\tau})$, let $J=\{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}:\text{$w_{\tau}=\mathrm{id},w_{3},w_{4},w_{5}$}\}$, and $J^{\prime}=\{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}:\text{$w_{\tau}\neq\mathrm{id},w_{3},w_{4},w_{5}$}\}$. For some choice of $((t_{\tau 1},t_{\tau 2})_{\tau},t)$ we see the equation (16) becomes
(20) $\prod_{\tau\in J^{\prime}}\chi_{\tau 1}\chi_{\tau 2}\prod_{\tau\in J}\chi_{\tau 1}^{-1}\simeq\mathbf{1}$
or
(21) $\prod_{\tau\in J^{\prime}}\chi_{\tau 1}\chi_{\tau 2}\prod_{\tau\in J}\chi_{\tau 1}^{-1}\simeq|\cdot|$
which contradicts genericity. Indeed, if we chose some $((t_{\tau 1},t_{\tau 2})_{\tau},t)$ for each $w_{\tau}$ such that we derived a contradiction in the case of $\mathrm{GSp}_{4}$ as above, then we see that the right-hand side of the isomorphism (16) simplifies to $|\cdot|^{n}$, for some $n$. If $n\neq 0,1$, then we see that by changing the values of $(t_{\tau 1},t_{\tau 2})$ for some $\tau\in J$, to the other choice of substitution, the right-hand side evaluates to $|\cdot|^{n-2}$, and continuing this process as necessary we get either equation (20) or (21).
@@
defines a minuscule highest weight representation, which, under the isomorphism $\mathrm{GSpin}_{5}\simeq\mathrm{GSp}_{4}$, identifies with the defining representation of $\mathrm{GSp}_{4}$. From here it is easy to see that the differences of the weights are roots of $\mathrm{GSp}_{4}$ (= coroots of $\mathrm{GSpin}_{5}$). For example, by using the parametrization of the maximal torus, as in (17), and the description of the roots in this parametrization provided on [Tad, Page 167]. Therefore, genericity guarantees strong $\mu$-regularity for this representation which implies $\mu$-regularity as before. The other fundamental tilting module of $\mathrm{GSpin}_{5}$ is given by the defining representation $\mathrm{GSpin}_{5}\rightarrow\mathrm{SO}_{5}\rightarrow\mathrm{GL}_{5}$ assuming $\ell\neq 2$ (See [Ham2, Appendix B]). Moreover, this occurs as a $5$-dimensional summand of $V_{\mathrm{spin}}\otimes V_{\mathrm{spin}}$, and it follows by [Ham2, Proposition 10.12] that we know $\mu$-regularity for this representation as well. Therefore, since we know $\mu$-regularity for the fundamental coweights, we are now done by [Ham2, Corollary 10.13] as before.
Now, for case (2), note that, as in the previous lemma, all cocharacters of $H$ are of the form $\mu=\prod_{\tau}\mu_{\tau}$ for some cocharacters $\mu_{\tau}$ of $\mathrm{GSp}_{4}$. The argument given above for $\mathrm{GSp}_{4}$ shows that if we let $\mu^{\prime}$ be one of the fundamental coweights of $\mathrm{GSp}_{4}$, then if we take $\mu=\prod\mu^{\prime}$, (i.e. $\mu_{\tau}=\mu^{\prime}$ for all $\tau$) then we are $\mu$-regular for such $\mu$. Applying [Ham2, Corollary 10.13] again shows that if $\mu=\prod_{\tau}\mu^{\prime}$, where $\mu^{\prime}$ is a dominant cocharacter of $\mathrm{GSp}_{4}$, then $\phi_{T}$ will be $\mu$-regular for such $\mu$. Note that, as in the case of $G(\mathrm{SL}_{n})$, every dominant minuscule cocharacter of $H$ is of this form. ∎
Now let us package our final result in a nice form. We first consider the following table, summarizing the groups and primes for which our results apply. We have left the entry blank if no constraint is imposed, and just mentioned the groups that appear as local constituents of global groups that admit a Shimura datum and for which $G$ is unramified.
[…]
@@
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ $L=\mathbb{Q}_{p}$ $(\ell,2(p^{4}-1))=1$
$L/\mathbb{Q}_{p}$ unramified $(\ell,2[L:\mathbb{Q}_{p}](p^{4[L:\mathbb{Q}_{p}]}-1))=1$ $p\neq 2$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$
@@
$G(\mathrm{Sp}_{4,L})$ $L/\mathbb{Q}_{p}$ unramified, $L\neq\mathbb{Q}_{p}$ $(\ell,2[L:\mathbb{Q}_{p}](p^{4[L:\mathbb{Q}_{p}]}-1))=1$ $p\neq 2$
@@
Assume $G$ is a product of the groups appearing in Table (22) with $p$ and $\ell$ satisfying the corresponding conditions. Then, we have that the natural map
$j_{1}^{*}T_{\mu}(-):\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\rightarrow\mathrm{D}^{\mathrm{adm}}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$
is exact with respect to the perverse $t$-structure on the source and the natural $t$-structure (= perverse $t$-structure) on the target for all minuscule $\mu$.
@@
Observe that all the groups in Table (22) satisfy Assumption 4.4, where the first five rows follows from Theorem 4.10, and the last two from Proposition 4.8. We now apply Theorem 4.20. To do this, we also need to check that if $\phi_{T}$ is a generic toral parameter then is also weakly normalized regular, and $\mu$-regular for all minuscule $\mu$. This follows from Lemma 4.22 and Lemma 4.23.
@@
Lastly, we need to check that the representations $\rho_{b,w}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ are semi-simple for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$. We claim that they are in fact irreducible. Recall that $J_{b}\simeq M_{b}\subset G$, where $M_{b}$ is a Levi of $G$. Moreover, we note that any such Levi $M_{b}$ is a product of groups also appearing in (22). Therefore, the desired irreducibility follows from the $\mu$-regularity, weak normalized regularity, and regularity of $\phi_{T}$ combined with Lemma 4.18. Note that in the case of $G(\mathrm{SL}_{2,L})$ and $G(\mathrm{Sp}_{4,L})$, we can always find a cocharacter $\mu$ of the form $\prod_{\tau}\mu^{\prime}$ which is not fixed by the Weyl group, since we can simply look at any cocharacter of $\mu^{\prime}$ of $\mathrm{GL}_{2}$ (resp. $\mathrm{GSp}_{4}$) which is not fixed by the Weyl group, and take $\mu$ to be the product of these $\mu^{\prime}$. ∎
@@
Assume $G$ is a product of the groups appearing in Table (22) with $p$ and $\ell$ satisfying the corresponding conditions. Then, for $\mathfrak{m}\subset H_{K_{p}}^{\mathrm{hs}}$ a generic maximal ideal, we have that
$\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\simeq\bigoplus_{b\in B(G)_{\mathrm{un}}}\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}.$
Moreover, the !!! and * pushforwards agree for any $A\in\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\simeq\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$
@@
5. The Proof of Theorems 1.14 and 1.16

### v5

@@ 4.4.2. Groups of type C
Thus, if we wanted to argue by contradiction, using the notation of the proof above, when $w_{\tau}=w_{i}$ for $i=1,\dots,7$, we should substitute for $t_{\tau 1},t_{\tau 2},t$ the values we considered above, subject to the additional constraint that we must have $t$, the similitude factor, being equal for all $\tau$. However, in all of the above cases $t=x$, and we choose different values for $t_{1},t_{2}$ depending on the particular $w_{i}$. In particular, as in the previous proof, to derive a contradiction we just need to evaluate the relationship
$\chi\simeq\chi^{w}$
on $t=x$ and set $t_{\tau 1}=1$ or $x$ and $t_{\tau 2}$ or $x$ depending on $w_{\tau}$, as in the proof described above. Since the roots of $H$ are expressible as products of roots of $\mathrm{GSp}_{4}$, this will allow us to derive a contradiction to genericity by the same argument as above.
@@
defines a minuscule highest weight representation, which, under the isomorphism $\mathrm{GSpin}_{5}\simeq\mathrm{GSp}_{4}$, identifies with the defining representation of $\mathrm{GSp}_{4}$. From here it is easy to see that the differences of the weights are roots of $\mathrm{GSp}_{4}$ (= coroots of $\mathrm{GSpin}_{5}$). For example, by using the parametrization of the maximal torus, as in (21), and the description of the roots in this parametrization provided on [63, Page 167]. Therefore, genericity guarantees strong $\mu$-regularity for this representation which implies $\mu$-regularity as before. The other fundamental tilting module of $\mathrm{GSpin}_{5}$ is given by the defining representation $\mathrm{GSpin}_{5}\rightarrow\mathrm{SO}_{5}\rightarrow\mathrm{GL}_{5}$ or its maximal irreducible submodule. Moreover, this occurs as a $5$-dimensional summand of $V_{\mathrm{spin}}\otimes V_{\mathrm{spin}}$, and it follows by [24, Proposition 9.12] that we know $\mu$-regularity for this representation as well. Therefore, since we know $\mu$-regularity for the fundamental coweights, we are now done by [24, Corollary 9.13] as before.
Now, for case (2), note that, as in the previous Lemma, all cocharacters of $H$ are of the form $\mu=\prod_{\tau}\mu_{\tau}$ for some cocharacters $\mu_{\tau}$ of $\mathrm{GSp}_{4}$. The argument given above for $\mathrm{GSp}_{4}$ shows that if we let $\mu^{\prime}$ be one of the fundamental coweights of $\mathrm{GSp}_{4}$, then if we take $\mu=\prod\mu^{\prime}$, (i.e. $\mu_{\tau}=\mu^{\prime}$ for all $\tau$) then we are $\mu$-regular for such $\mu$. Applying [24, Corollary 9.13] again shows that if $\mu=\prod_{\tau}\mu^{\prime}$, where $\mu^{\prime}$ is a dominant cocharacter of $\mathrm{GSp}_{4}$, then $\phi_{T}$ will be $\mu$-regular for such $\mu$. Note that, as in the case of $G(\mathrm{SL}_{2})$, every dominant minuscule cocharacter of $H$ is of this form. ∎
With all the extra conditions now properly understood in the relevant cases, we now turn our attention to formulating and proving the main local results.
4.4.3. The Local Results
We consider the following table, summarizing the groups and primes for which our results apply. We have left the entry blank if no constraint is imposed, and just mentioned the groups that appear as local constituents of global groups that admit a Shimura datum and for which $G$ is unramified.
$G$ Constraint on $G$ $\ell$ $p$
@@ 4.4.3. The Local Results
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ $L=\mathbb{Q}_{p}$
$L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$ $p\neq 2$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ $L/\mathbb{Q}_{p}$ unramified, $(\ell,2[L:\mathbb{Q}_{p}])=1$
@@
$G(\mathrm{Sp}_{4,L})$ $L/\mathbb{Q}_{p}$ unramified, $L\neq\mathbb{Q}_{p}$ $(\ell,[L:\mathbb{Q}_{p}])=1$ $p\neq 2$
@@
Assume $G$ is a product of the groups appearing in Table (23) with $p$ and $\ell$ satisfying the corresponding conditions. Let $\phi$ be a semisimple $L$-parameter which is induced from a generic toral parameter $\phi_{T}$. Let $\mu$ be a geometric dominant cocharacter of $G$ which satisfies the constraint that it is of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$, for $\mu^{\prime}$ a cocharacter of $\mathrm{GL}_{2}$ (resp. $\mathrm{GSp}_{4}$) on the direct factors of $G$ equal to $G(\mathrm{SL}_{2,L})$ or $G(\mathrm{Sp}_{4,L})$, then we have that the natural functor
$T_{\mu}:\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\rightarrow\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
is perverse t-exact. In particular, it is perverse $t$-exact for all minuscule $\mu$, and if we look at the natural functor
$j_{1}^{*}T_{\mu}:\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\rightarrow\mathrm{D}^{\mathrm{adm}}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$
then this is exact for the perverse $t$-structure on the source and the perverse (= standard $t$-structure) on the target.
@@
Observe that all the groups in Table (23) satisfy Assumption 4.5, where the first five rows follows from Theorem 4.13, and the last two from Proposition 4.9. We also note that the $\ell$ very decent assumption is also guaranteed by our assumptions on $\ell$ as in Theorem 4.13, so we may use the results of §4.3. We now apply Theorem 4.23. To do this, we also need to check that if $\phi_{T}$ is a generic toral parameter then it is also generic regular, and $\mu$-regular for all $\mu$ if $G\neq\mathrm{Res}_{L/\mathbb{Q}_{p}}H$ and $H=\mathrm{GL}_{n},\mathrm{GU}_{n}$, or $\mathrm{U}_{n}$. If $G=G(\mathrm{SL}_{2,L})$ (resp. $G(\mathrm{Sp}_{4,L})$), and we need to check if $\phi_{T}$ is generic then it is generic regular and $\mu$-regular for all $\mu$ of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$, for $\mu^{\prime}$ a cocharacter of $\mathrm{GL}_{2}$ (resp. $\mathrm{GSp}_{4}$) it is $\mu$-regular, where these give rise to all minuscule cocharacters of $G$ in this case. Finally, if $G=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}$ we need to check if $\phi$ is generic then it is regular and $\mu$-regular for all $\mu$. These claims follow from Lemma 4.25 and Lemma 4.26.
@@
Lastly, we need to check that the representations $\rho_{b,w}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ are semi-simple for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$. We claim that they are in fact irreducible. Recall that $J_{b}\simeq M_{b}\subset G$, where $M_{b}$ is a Levi of $G$. Moreover, we note that any such Levi $M_{b}$ is a product of groups also appearing in (23). Therefore, the desired irreducibility follows from the $\mu$-regularity and regularity of $\phi_{T}$ discussed above combined with Lemma 4.21. Note that in the case of $G(\mathrm{SL}_{2,L})$ and $G(\mathrm{Sp}_{4,L})$, we can always find a cocharacter $\mu$ of the form $\prod_{\tau}\mu^{\prime}$ which is not fixed by the Weyl group, since we can simply look at any cocharacter of $\mu^{\prime}$ of $\mathrm{GL}_{2}$ (resp. $\mathrm{GSp}_{4}$) which is not fixed by the Weyl group, and take $\mu$ to be the product of these $\mu^{\prime}$ in order to see that the assumptions of Lemma 4.21 are satisfied. ∎
Remark 4.28.
We note that if the local group $G$ is split then the set $B(G,\mu)_{\mathrm{un}}$ is a singleton for any minuscule cocharacter $\mu$, by invoking [24, Corollary 2.9]. We denote the unique element in this set as $b_{\mu}$. In this case, for a semisimple $L$-parameter $\phi$ induced from a generic $\phi_{T}$, by using Proposition 4.6 and that assumption 4.5 holds, we see that the adjunction map $j_{b_{\mu}!}j_{b_{\mu}}^{!}\rightarrow\mathrm{id}$ induces a natural isomorphism
$j_{1}^{*}T_{\mu}j_{b_{\mu}!}j_{b_{\mu}}^{*}\simeq j_{1}^{*}T_{\mu}$
of operators on the full subcategory $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G})$ (Here we have implicitly used the isomorphism $j_{1}^{*}T_{\mu}\simeq j_{1}^{*}T_{\mu}j_{\leq\mu!}j_{\leq\mu}^{*}$, by [53, Proposition A.9] , to rewrite $j_{b_{\mu}}^{!}$ as $j_{b_{\mu}}^{*}$ using Proposition 4.6, where $j_{\leq\mu}:\mathrm{Bun}_{G,\leq\mu}\hookrightarrow\mathrm{Bun}_{G}$ is the open immersion of the substack corresponding to $B(G,\mu)\subset B(G)$). We have that the $\sigma$-centralizer of $J_{b_{\mu}}$ identifies with a proper Levi of $M_{b_{\mu}}\subset G$ and the LHS of this isomorphism can be explicitly computed (up to a shift by the dimension of the Shimura variety) in terms of the parabolic induction from $J_{b_{\mu}}\simeq M_{b_{\mu}}$ to $G$, using the results of [18]. In this case, the second part of Corollary 4.27 is a simple consequence of the exactness of the parabolic induction functor (even without restricting to the ULA subcategory). In particular, for any of cases appearing in Table (23) where the group $G$ is split, Corollary 4.27 also holds immediately upon knowing the that an object in $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is supported on just the unramified elements for $\phi$ induced from a generic toral parameter $\phi_{T}$, which follows from Assumption 4.5 and Proposition 4.6. As described in [37, 54] in the case when $G$ is a product of $\mathrm{GL}_{n}$s, by an observation of Koshikawa one can actually weaken the condition on $\phi$ to just assuming that $H^{2}(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})\simeq 0$ (cf. Remark 6.8) to deduce that $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ is supported on the unramified elements from assumption 4.5. The key point here is that non-trivial classes in the $H^{0}(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ will only give rise to non Frobenius semi-simple deformations of $\phi_{T}$ as a $\phantom{}{}^{L}T(\overline{\mathbb{F}}_{\ell})$-valued parameter to a $\phantom{}{}^{L}B(\overline{\mathbb{F}}_{\ell})$-valued parameter, which cannot occur as mod $\ell$-reductions of the semi-simplification of parameters appearing in the classical Local Langlands correspondence (cf. Remark 6.5). By the reasoning described above, this will in turn imply Conjecture 1.2 under this weaker genericity condition in the case when $G$ is split (See [10, Proposition 10.2.4,10.2.5] for details). However, under this weaker hypothesis, the splitting of the semi-orthogonal decomposition (See Corollary 4.29 below) should not hold (cf. Remark 6.8) and this becomes important when $B(G,\mu)_{\mathrm{un}}$ has more than one element.
@@
Assume $G$ is a product of the groups appearing in Table (23) with $p$ and $\ell$ satisfying the corresponding conditions. Then, for $\phi$ a semisimple parameter induced from a generic $\phi_{T}$, we have that
$\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\bigoplus_{b\in B(G)_{\mathrm{un}}}\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p},\overline{\mathbb{F}}_{\ell})_{\phi}.$
Moreover, the !! and * pushforwards agree for any $A\in\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}$
@@
5. The Proof of the Main Theorems
We recall that the pair $(\mathbf{G},X)$ will denote a Shimura datum with reflex field $E/\mathbb{Q}$. We let $\ell\neq p$ be distinct prime numbers, and $E_{\mathfrak{p}}$ the completion at the place $\mathfrak{p}$ in $E$ dividing $p$ determined by the fixed isomorphism $j:\overline{\mathbb{Q}}_{p}\xrightarrow{\simeq}\mathbb{C}$ with completed algebraic closure denoted by $C$. We will let $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ be the local group with maximal torus $T$, and we let $\mu$ be the conjugacy class of geometric dominant cocharacters of $G$ attached to the inverse of the Hodge cocharacter of $X$ and the isomorphism $j:\overline{\mathbb{Q}}_{p}\xrightarrow{\simeq}\mathbb{C}$. We assume that $G$ is unramified, and let $K_{p}^{\mathrm{hs}}\subset G(\mathbb{Q}_{p})$ denote a hyperspecial level. We let $H_{K_{p}^{\mathrm{hs}}}$ denote the spherical Hecke algebra with coefficients in $\overline{\mathbb{F}}_{\ell}$, and recall that $B(G)_{\mathrm{un}}:=\mathrm{Im}(B(T)\rightarrow B(G))$ denotes the set of unramified elements.
For $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we write $\mathrm{Sh}(\mathbf{G},X)_{K}/\mathop{\rm Spec}(E)$ for the Shimura variety attatched to the Shimura datum, with its associated adic space $\mathcal{S}(\mathbf{G},X)_{K}$ over $\mathop{\rm Spa}(E_{\mathfrak{p}})$. and if $K=K^{p}K_{p}$, where $K_{p}$ (resp. $K^{p}$) denotes the level at $p$ (resp. away from $p$), we write $\mathcal{S}(\mathbf{G},X)_{K^{p}}$ for the associated Shimura variety at infinite level introduced in §3.1.
In this section, we will reap the fruit of our work in the previous section and finally prove our main results. More precisely, in §5.1, we will establish our main results on the splitting of Mantovan’s filtration (Theorem 5.1) for the torsion cohomology of a PEL type AC Shimura variety, as well as show the bounds for the torsion cohomology after generic localization (Theorem 5.2). In §5.2, we push our results a bit further and deduce additional applications to abelian type Shimura varieties (Corollaries 5.6, 5.7).